# On the non-Markovian quantum control dynamics

### v2

@@ IV-B Feedback control with cavity operators
$\langle\dot{x}\rangle=\Delta\langle p\rangle-\frac{\kappa}{2}\langle x\rangle-i\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle+i\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}+\sqrt{2}g_{f}\beta_{p}\langle x\rangle+\sqrt{2}\left(V_{x}-\frac{1}{2}\right)\xi(t)+g_{f}\beta_{p}\xi(t),$ (63a)
$\langle\dot{p}\rangle=-\Delta\langle x\rangle-\frac{\kappa}{2}\langle p\rangle-\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle-\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}-\sqrt{2}g_{f}\beta_{x}\langle x\rangle+\sqrt{2}V_{xp}\xi(t)-g_{f}\beta_{x}\xi(t),$ (63b)
$\langle\dot{\sigma}^{-}_{n}\rangle=-ig_{n}e^{i\delta t}\langle a\rangle-F(t)|\kappa_{n}|^{2}\langle\sigma^{-}_{n}\rangle,$ (63c)
$\langle\dot{a}\rangle=-i\Delta\langle a\rangle-i\sum_{n}g_{n}e^{-i\delta t}\langle\sigma^{-}_{n}\rangle-\frac{\kappa}{2}\langle a\rangle-ig_{f}\left(\beta_{x}+i\beta_{p}\right)\langle x\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}+\left(\left\langle\left(a+a^{{\dagger}}\right)a\right\rangle-\left\langle a+a^{{\dagger}}\right\rangle\left\langle a\right\rangle\right)\xi(t)-\frac{ig_{f}}{\sqrt{2}}\left(\beta_{x}+i\beta_{p}\right)\xi(t),$ (63d)
where $V_{x}$ is the variance of the position operator as $V_{x}=\left\langle x^{2}\right\rangle-\langle x\rangle^{2}$, $V_{xp}$ is the covariance between position and momentum as $V_{xp}=\left(\langle xp\rangle+\langle px\rangle\right)/2-\langle x\rangle\langle p\rangle$ [77], and $F(t)$ is governed by Eq. (9e).
@@
$\dot{V}_{x}=2\Delta V_{xp}-\kappa V_{x}^{\prime}+2\sqrt{2}g_{f}\beta_{p}V_{x}^{\prime}+g_{f}^{2}\beta_{p}^{2}-\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right)^{2},$ (64a)
$\dot{V}_{xp}=\Delta V_{p}-\Delta V_{x}-\kappa V_{xp}+\sqrt{2}g_{f}\beta_{p}V_{xp}-\sqrt{2}g_{f}\beta_{x}V_{x}$
$~{}~{}~{}~{}~{}~{}~{}~{}+\frac{\sqrt{2}}{2}g_{f}\beta_{x}-g_{f}^{2}\beta_{x}\beta_{p}-\left(\sqrt{2}V_{xp}-g_{f}\beta_{x}\right)\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right),$ (64b)
@@
For the open loop control without feedback, Eq. (64) can finally converge to the steady values $V_{x}(\infty)=V_{p}(\infty)=\frac{1}{2}$ and $V_{xp}(\infty)=0$, similar as in Ref. [81].
For arbitrary $\beta_{x}$, $\beta_{p}$ and $g_{f}$, Eq. (64) can be simplified as
$\dot{V}_{x}=2\Delta V_{xp}-\kappa V_{x}^{\prime}-2V_{x}^{\prime 2},$ (65a)
$\dot{V}_{xp}=\Delta V_{p}-\Delta V_{x}-\kappa V_{xp}-2V_{x}^{\prime}V_{xp},$ (65b)
$\dot{V}_{p}=-2\Delta V_{xp}-\kappa V_{P}^{\prime}-2V_{xp}^{2},$ (65c)
where $V_{p}^{\prime}=V_{p}-1/2$ similar as the definition of $V_{x}^{\prime}$. It can be seen that the steady values for valience can be influenced by detuning $\Delta$ of the feedback field. Obviously, when $\Delta=0$, $V_{xp}(\infty)=0$, $V_{x}(\infty)=V_{p}(\infty)=1/2$, which reduces to the circumstance in Ref. [81].
@@ V Quantum control for the multiple Jaynes-Cummings model
When there are overall $M$ nearest neighbor coupled cavities with one two-level atom in each cavity, the free Hamiltonian for the atoms and cavities reads
$\bar{H}_{M}$ $=\sum_{m=1}^{M}\omega_{c}^{(m)}a^{{\dagger}}_{m}a_{m}+\sum_{m=1}^{M}\omega_{a}^{(m)}\sigma_{+}^{(m)}\sigma_{-}^{(m)},$ (66)
@@
where $g_{m}$ is the coupling strength between the atom and cavity in the $m$-th cavity, $J_{c}$ is the coupling between two neighborhood cavities, the atom in each cavity is coupled with the environment via non-Markovian dynamics as
$H_{A-E}$ $=\sum_{m=1}^{M}\sum_{\omega}\chi_{\omega}^{(m)}\left(\sigma_{+}^{(m)}b_{\omega}+\sigma_{-}^{(m)}b_{\omega}^{{\dagger}}\right),$ (68)
and the atomic system is driven by the non-Markovian master equation as
$\mathrm{d}\rho=$$\left\{-i\left[H_{M},\rho(t)\right]+\sum_{m}\left[\left(F_{m}^{*}(t)+F_{m}(t)\right)\sigma_{-}^{(m)}\rho(t)\sigma_{+}^{(m)}\right.\right.$ (69)
$\left.\left.-F_{m}(t)\sigma_{+}^{(m)}\sigma_{-}^{(m)}\rho(t)-F_{m}^{*}(t)\rho(t)\sigma_{+}^{(m)}\sigma_{-}^{(m)}+\kappa_{m}\mathcal{L}_{a_{m}}\left[\rho(t)\right]\right]\right\}\mathrm{d}t,$
where $\sigma_{-}^{(m)}=I_{1}\otimes\cdots\otimes I_{m-1}\otimes\sigma_{-}\otimes I_{m+1}\cdots\otimes I_{M}$ and
$\frac{\mathrm{d}}{\mathrm{d}t}F_{m}(t)=\left|\kappa^{(m)}\right|^{2}F_{m}^{2}(t)-\left(\gamma+i\Omega-i\omega_{a}^{(m)}\right)F_{m}(t)+\frac{\gamma\chi}{2},$
where $\kappa^{(m)}$ represents the coupling between the atom in the $m$-th cavity and environment.
@@ V-A Dynamics without drives
When there are no drives applied either to the cavities or atoms, generalized from Eq. (9), the open-loop dynamics in the $m$-th cavity reads [78, 82]
$\frac{\mathrm{d}}{\mathrm{d}t}\langle a_{m}^{{\dagger}}a_{m}\rangle=$ $ig_{m}\left(e^{i\delta t}\langle\sigma_{+}^{(m)}a_{m}\rangle-e^{-i\delta t}\langle\sigma_{-}^{(m)}a^{{\dagger}}_{m}\rangle\right)-\kappa\langle a_{m}^{{\dagger}}a_{m}\rangle$ (70)
$-iJ_{c}\langle a_{m}^{{\dagger}}a_{m+1}\rangle+iJ_{c}\langle a_{m}a_{m+1}^{{\dagger}}\rangle+iJ_{c}\langle a_{m-1}^{{\dagger}}a_{m}\rangle-iJ_{c}\langle a_{m-1}a_{m}^{{\dagger}}\rangle,$
$\frac{\mathrm{d}}{\mathrm{d}t}\langle a_{m}a_{m+1}^{{\dagger}}\rangle=iJ_{c}\left(\langle a_{m}^{{\dagger}}a_{m}\rangle-\langle a_{m+1}^{{\dagger}}a_{m+1}\rangle\right).$ (71)
Besides, the dynamics of the operators in the $m$-th cavity $\langle\sigma_{+}^{(m)}a_{m}\rangle$ and $\langle\sigma_{-}^{(m)}a_{m}^{{\dagger}}\rangle$ are the same as that of $\langle\sigma^{+}_{n}a\rangle$ and $\langle\sigma^{-}_{n}a^{{\dagger}}\rangle$ in Eq. (9) respectively by taking $n=N-1=1$ and omitting the higher order components.
Denote $\mathbf{x}_{m}=\left[\langle\sigma_{-}^{(m)}\rangle,\langle a_{m}\rangle\right]^{T}$, and $\tilde{\mathbf{X}}=\left[\mathbf{x}_{1}^{T},\mathbf{x}_{2}^{T},\cdots,\mathbf{x}_{M}^{T}\right]^{T}$, then
$\dot{\mathbf{x}}_{m}=$ $\begin{bmatrix}-F_{m}(t)|\kappa^{(m)}|^{2}&-ig_{m}e^{i\delta_{m}t}\\ -ig_{m}e^{-i\delta_{m}t}&-\kappa\end{bmatrix}\mathbf{x}_{m}$ (72)
$+\begin{bmatrix}0\\ -iJ_{c}\end{bmatrix}\langle a_{m+1}\rangle+\begin{bmatrix}0\\ -iJ_{c}\end{bmatrix}\langle a_{m-1}\rangle.$
V-A1 Multi-cavity case
@@
$\mathbf{\mathcal{X}}_{M}=\left[\mathbf{s}^{R}_{1},\mathbf{s}^{I}_{1},\cdots,\mathbf{s}^{R}_{M},\mathbf{s}^{I}_{M},\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\cdots,\mathbf{a}^{R}_{M},\mathbf{a}^{I}_{M}\right]^{T},$
@@
$\mathbf{F}(t)=\begin{bmatrix}-\mathbf{\mathcal{A}}^{F}_{1}(t)|\kappa^{(1)}|^{2}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&-\mathbf{\mathcal{A}}^{F}_{M}(t)|\kappa^{(M)}|^{2}\end{bmatrix},$ (74)
$\mathbf{G}(t)=\begin{bmatrix}\mathbf{\mathcal{A}}^{G}_{1}(t)&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&\mathbf{\mathcal{A}}^{G}_{M}(t)\end{bmatrix},$ (75)
$\mathbf{R}(t)=\begin{bmatrix}\mathbf{\mathcal{A}}^{g}_{1}(t)&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&\mathbf{\mathcal{A}}^{g}_{M}(t)\end{bmatrix},$ (76)
$\mathbf{S}$ $=\begin{bmatrix}\mathbf{\mathcal{A}}_{\kappa}&\mathbf{\mathcal{A}}_{J}&0&\cdots&0\\ \mathbf{\mathcal{A}}_{J}&\mathbf{\mathcal{A}}_{\kappa}&\mathbf{\mathcal{A}}_{J}&\cdots&0\\ 0&\mathbf{\mathcal{A}}_{J}&\mathbf{\mathcal{A}}_{\kappa}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\mathbf{\mathcal{A}}_{\kappa}\end{bmatrix}$ (77)
$=I_{M}\otimes\mathbf{\mathcal{A}}_{\kappa}+\begin{bmatrix}0&1&0&\cdots&0\\ 1&0&1&\cdots&0\\ 0&1&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0\end{bmatrix}\otimes\mathbf{\mathcal{A}}_{J},$
where $\mathbf{\mathcal{A}}^{F}_{j}(t)=\begin{bmatrix}F^{R}_{j}(t)&-F^{I}_{j}(t)\\ F^{I}_{j}(t)&F^{R}_{j}(t)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{G}_{j}(t)=\begin{bmatrix}g_{j}\sin\left(\delta_{j}t\right)&g_{j}\cos\left(\delta_{j}t\right)\\ -g_{j}\cos\left(\delta_{j}t\right)&g_{j}\sin\left(\delta_{j}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{g}_{j}(t)=\begin{bmatrix}-g_{j}\sin\left(\delta_{j}t\right)&g_{j}\cos\left(\delta_{j}t\right)\\ -g_{j}\cos\left(\delta_{j}t\right)&-g_{j}\sin\left(\delta_{j}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}_{\kappa}=\begin{bmatrix}-\kappa&0\\ 0&-\kappa\end{bmatrix}$ and $\mathbf{\mathcal{A}}_{J}=\begin{bmatrix}0&J_{c}\\ -J_{c}&0\end{bmatrix}$.
@@
When the interaction between the atom and the environment is Markovian or converges from non-Markovian to Markovian interactions, $F_{j}(t)$ are constants and $\mathbf{\mathcal{A}}_{M}$ in Eq. (73) is periodic as $\mathbf{\mathcal{A}}_{M}\left(t+T_{M}\right)=\mathbf{\mathcal{A}}_{M}\left(t\right)$ and $T_{M}=\frac{2\pi}{\delta_{1}}$ when $\delta_{1}=\cdots=\delta_{M}\neq 0$. Further when $\delta_{1}=\cdots=\delta_{M}=0$, $\mathbf{\mathcal{A}}_{M}\left(t\right)$ converges to be time invariant.

### v3

@@ IV-A Feedback control with cavity operators
$\langle\dot{x}\rangle=\Delta\langle p\rangle-\frac{\kappa}{2}\langle x\rangle$
$~{}~{}~{}~{}~{}~{}~{}-i\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle+i\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}+\sqrt{2}g_{f}\beta_{p}\langle x\rangle+\sqrt{2}\left(V_{x}-\frac{1}{2}\right)\xi(t)+g_{f}\beta_{p}\xi(t),$ (55a)
$\langle\dot{p}\rangle=-\Delta\langle x\rangle-\frac{\kappa}{2}\langle p\rangle$
$~{}~{}~{}~{}~{}~{}~{}-\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle-\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}-\sqrt{2}g_{f}\beta_{x}\langle x\rangle+\sqrt{2}V_{xp}\xi(t)-g_{f}\beta_{x}\xi(t),$ (55b)
$\langle\dot{\sigma}^{-}_{n}\rangle=-ig_{n}e^{i\delta t}\langle a\rangle-F_{n}(t)\kappa_{n}\langle\sigma^{-}_{n}\rangle,$ (55c)
$\langle\dot{a}\rangle=-i\Delta\langle a\rangle-i\sum_{n}g_{n}e^{-i\delta t}\langle\sigma^{-}_{n}\rangle-\frac{\kappa}{2}\langle a\rangle$
$~{}~{}~{}~{}~{}~{}~{}-ig_{f}\left(\beta_{x}+i\beta_{p}\right)\langle x\rangle-\frac{ig_{f}}{\sqrt{2}}\left(\beta_{x}+i\beta_{p}\right)\xi(t)$
$~{}~{}~{}~{}~{}~{}~{}+\left[\left\langle\left(a+a^{{\dagger}}\right)a\right\rangle-\left\langle a+a^{{\dagger}}\right\rangle\left\langle a\right\rangle\right]\xi(t),$ (55d)
where $V_{x}$ is the variance of the position operator as $V_{x}=\left\langle x^{2}\right\rangle-\langle x\rangle^{2}$, $V_{xp}$ is the covariance between position and momentum as $V_{xp}=\left(\langle xp\rangle+\langle px\rangle\right)/2-\langle x\rangle\langle p\rangle$ [63], and $F_{n}(t)$ is governed by Eq. (9e).
@@
$\dot{V}_{x}=2\Delta V_{xp}-\kappa V_{x}^{\prime}+2\sqrt{2}g_{f}\beta_{p}V_{x}^{\prime}+g_{f}^{2}\beta_{p}^{2}$
$~{}~{}~{}~{}~{}~{}~{}-\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right)^{2},$ (56a)
$\dot{V}_{xp}=\Delta V_{p}-\Delta V_{x}-\kappa V_{xp}+\sqrt{2}g_{f}\beta_{p}V_{xp}$
$~{}~{}~{}~{}~{}~{}~{}~{}-\sqrt{2}g_{f}\beta_{x}V_{x}+\frac{\sqrt{2}}{2}g_{f}\beta_{x}-g_{f}^{2}\beta_{x}\beta_{p}$
$~{}~{}~{}~{}~{}~{}~{}~{}-\left(\sqrt{2}V_{xp}-g_{f}\beta_{x}\right)\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right),$ (56b)
@@
For the open loop control without feedback, the variances in Eq. (56) converge to the steady values as $V_{x}(\infty)=V_{p}(\infty)=0.5$ and $V_{xp}(\infty)=0$, similar as in Ref. [67].
IV-B Feedback control with atomic operators
The feedback operator for the $N$-level system can be represented as [15]
When $N=2$, we choose $G=|0\rangle\langle 1|+|1\rangle\langle 0|$, and assume initially the atom is excited. The system’s states can be evaluated with the vector of the operators $S=\left[\langle a\rangle,\langle\sigma_{-}\rangle,\langle\sigma_{z}\rangle\right]^{\rm T}\triangleq\left[\bar{\alpha},\bar{s},\bar{w}\right]^{\rm T}$, where $\sigma_{-}=|0\rangle\langle 1|$ and $\sigma_{z}=|1\rangle\langle 1|-|0\rangle\langle 0|$. When the external drive is $H_{D}=-i\mathcal{E}\left(a-a^{{\dagger}}\right)$, the dynamics of the cavity or atomic operator reads
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{\alpha}=-ig_{1}e^{-i\delta t}\bar{s}-\frac{F_{a}(t)\kappa}{2}\bar{\alpha}+\mathcal{E},$ (58a)
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{s}=ig_{1}e^{i\delta t}\bar{\alpha}\bar{w}-F_{1}(t)\kappa_{1}\bar{s}+ig_{f}\left(\bar{\alpha}+\bar{\alpha}^{*}\right)\bar{w}$
$-\left(F_{1}(t)+F_{1}^{*}(t)\right)\kappa_{1}\left(\bar{w}+1\right)-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\left(\bar{\alpha}+\bar{\alpha}^{*}\right)$
$-2g_{f}^{2}\bar{w}-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\xi(t),$ (58c)
with the initially normalization condition $\bar{w}^{2}+4\left|\bar{s}\right|^{2}=1$ [68], and the stochastic component with higher order amplitude is omitted in Eq. (58a). We consider the simplified steady states satisfying that $\dot{\bar{\alpha}}=\dot{\bar{s}}=\dot{\bar{w}}=0$ after averaging the measurement noise $\xi(t)=0$, and take $\delta=0$ for simplification [20, 69].
Figure 5: Measurement feedback control for one two-level atom in the cavity when both the atom and cavity interact with a non-Markovian environment.
The performance for different non-Markovian parameter settings is compared in Fig. 5, where we take $\kappa_{1}=1$GHz, $\kappa=0.2$GHz, $\gamma=2$GHz, $\chi=1$, $\omega_{a}=37.7$GHz, $\Omega=32.7$GHz, $g_{1}=200$MHz, $\delta=0$MHz and $\mathrm{d}t=0.01$ns. It can be seen that the measurement feedback control can enhance the probability that the atom is excited, especially when the feedback strength and drive field are strong.
@@ V Quantum control for the multiple Jaynes-Cummings model
Based on the above analysis on the non-Markovian quantum dynamics with one cavity, in this section, we further study the non-Markovian dynamics that atoms are in an array of coupled cavities. When there are overall $M$ nearest neighbor coupled cavities with one two-level atom in each cavity, the free Hamiltonian for the atoms and cavities reads
$H_{M}$ $=\sum_{m=1}^{M}\omega_{c}^{(m)}a^{{\dagger}}_{m}a_{m}+\sum_{m=1}^{M}\omega_{a}^{(m)}\sigma_{+}^{(m)}\sigma_{-}^{(m)},$ (59)
@@
where $g_{m}$ is the coupling strength between the atom and cavity in the $m$-th cavity, $J_{c}$ is the coupling strength between two neighborhood cavities, the atom in each cavity is coupled to the non-Markovian environment similar as in Appendix A, and the atomic system is driven by the non-Markovian master equation as
$\mathrm{d}\rho=\left\{-i\left[H_{M},\rho\right]+\sum_{m}\left[\left(\mathcal{F}_{m}^{*}+\mathcal{F}_{m}\right)\sigma_{-}^{(m)}\rho\sigma_{+}^{(m)}\right.\right.$ (61)
$\left.\left.-\mathcal{F}_{m}\sigma_{+}^{(m)}\sigma_{-}^{(m)}\rho-\mathcal{F}_{m}^{*}\rho\sigma_{+}^{(m)}\sigma_{-}^{(m)}+\kappa_{m}\mathcal{L}_{a_{m}}\left[\rho\right]\right]\right\}\mathrm{d}t,$
where $\sigma_{-}^{(m)}=I_{1}\otimes\cdots\otimes I_{m-1}\otimes\sigma_{-}\otimes I_{m+1}\cdots\otimes I_{M}$,
$\frac{\mathrm{d}\mathcal{F}_{m}}{\mathrm{d}t}=\kappa^{(m)}\mathcal{F}_{m}^{2}-\left(\gamma+i\Omega-i\omega_{a}^{(m)}\right)\mathcal{F}_{m}+\frac{\gamma\chi}{2},$
$\kappa^{(m)}$ represents the coupling between the atom in the $m$-th cavity and environment, $\gamma$ and $\chi$ are constants.
@@ V-A Dynamics without drives
Denote $\mathbf{x}_{m}=\left[\langle\sigma_{-}^{(m)}\rangle,\langle a_{m}\rangle\right]^{\rm T}$, and $\tilde{\mathbf{X}}=\left[\mathbf{x}_{1}^{\rm T},\mathbf{x}_{2}^{\rm T},\cdots,\mathbf{x}_{M}^{\rm T}\right]^{\rm T}$, then
$\dot{\mathbf{x}}_{m}=$ $\begin{bmatrix}-\mathcal{F}_{m}(t)\kappa^{(m)}&-ig_{m}e^{i\delta_{m}t}\\ -ig_{m}e^{-i\delta_{m}t}&-\kappa_{m}\end{bmatrix}\mathbf{x}_{m}$ (62)
$-\begin{bmatrix}0\\ iJ_{c}\end{bmatrix}\langle a_{m+1}\rangle-\begin{bmatrix}0\\ iJ_{c}\end{bmatrix}\langle a_{m-1}\rangle.$
In the following, we denote and $\mathcal{F}_{m}(t)=\mathcal{F}^{R}_{m}(t)+i\mathcal{F}^{I}_{m}(t)$.
@@
$\mathbf{\mathcal{X}}_{M}=\left[\mathbf{s}^{R}_{1},\mathbf{s}^{I}_{1},\cdots,\mathbf{s}^{R}_{M},\mathbf{s}^{I}_{M},\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\cdots,\mathbf{a}^{R}_{M},\mathbf{a}^{I}_{M}\right]^{\rm T},$
@@
$\mathbf{F}(t)=-{\rm diag}\left[\mathbf{\mathcal{A}}^{F}_{1}(t)\kappa^{(1)},\cdots,\mathbf{\mathcal{A}}^{F}_{M}(t)\kappa^{(M)}\right],$ (64)
$\mathbf{G}(t)={\rm diag}\left[\mathbf{\mathcal{A}}^{G}_{1}(t),\cdots,\mathbf{\mathcal{A}}^{G}_{M}(t)\right],$ (65)
$\mathbf{R}(t)={\rm diag}\left[\mathbf{\mathcal{A}}^{g}_{1}(t),\cdots,\mathbf{\mathcal{A}}^{g}_{M}(t)\right],$ (66)
$\mathbf{S}$ $=I_{M}\otimes\mathbf{\mathcal{A}}_{\kappa}+\begin{bmatrix}0&1&0&\cdots&0\\ 1&0&1&\cdots&0\\ 0&1&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0\end{bmatrix}\otimes\mathbf{\mathcal{A}}_{J},$ (67)
where $\mathbf{\mathcal{A}}^{F}_{m}(t)=\begin{bmatrix}\mathcal{F}^{R}_{m}(t)&-\mathcal{F}^{I}_{m}(t)\\ \mathcal{F}^{I}_{m}(t)&\mathcal{F}^{R}_{m}(t)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{G}_{m}(t)=\begin{bmatrix}g_{m}\sin\left(\delta_{m}t\right)&g_{m}\cos\left(\delta_{m}t\right)\\ -g_{m}\cos\left(\delta_{m}t\right)&g_{m}\sin\left(\delta_{m}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{g}_{m}(t)=\begin{bmatrix}-g_{m}\sin\left(\delta_{m}t\right)&g_{m}\cos\left(\delta_{m}t\right)\\ -g_{m}\cos\left(\delta_{m}t\right)&-g_{m}\sin\left(\delta_{m}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}_{\kappa}=\begin{bmatrix}-\kappa&0\\ 0&-\kappa\end{bmatrix}$ and $\mathbf{\mathcal{A}}_{J}=\begin{bmatrix}0&J_{c}\\ -J_{c}&0\end{bmatrix}$.
@@
When the interaction between the atom and the environment is Markovian or converges from non-Markovian to Markovian interactions, $\mathcal{F}_{m}$ equal constants and $\mathbf{\mathcal{A}}_{M}$ in Eq. (63) is periodic as $\mathbf{\mathcal{A}}_{M}\left(t+2\pi/\delta_{m}\right)=\mathbf{\mathcal{A}}_{M}\left(t\right)$ when $\delta_{1}=\cdots=\delta_{M}\neq 0$. Besides, when $\delta_{1}=\cdots=\delta_{M}=0$, $\mathbf{\mathcal{A}}_{M}\left(t\right)$ reduces to be time invariant.
Obviously, the stability of the high-dimensional equation (63), which is linear time-varying, can be similarly clarified by Corollary 2. For example, when there are two coupled cavities, the dynamics based on Markovian interaction between the quantum system and the environment has been analyzed in Ref. [70]. In this case, we denote $\mathbf{\mathcal{X}}_{\mathbf{a}}=\left[\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\mathbf{a}^{R}_{2},\mathbf{a}^{I}_{2}\right]^{\rm T}$, as a subspace of $\mathbf{\mathcal{X}}$. Obviously, when $g_{m}=0$ for $m=1,2$,
$\dot{\mathbf{\mathcal{X}}_{\mathbf{a}}}$ $=\begin{bmatrix}\mathbf{\mathcal{A}}_{\kappa}&\mathbf{\mathcal{A}}_{J}\\ \mathbf{\mathcal{A}}_{J}&\mathbf{\mathcal{A}}_{\kappa}\end{bmatrix}\mathbf{\mathcal{X}}_{\mathbf{a}}\triangleq\mathbf{\mathcal{A}}_{\mathbf{a}}\mathbf{\mathcal{X}}_{\mathbf{a}},$ (68)
which is linear and time-invariant. $\mathbf{\mathcal{X}}_{\mathbf{s}}$ is exponentially stable because in Eq. (68) the real part of roots determined by $(\textit{s}+\kappa)^{2}+J_{c}^{2}=0$ are negative, where s is for the Laplace transformation. Further by Corollary 1, $\lim_{t\rightarrow\infty}\mathbf{\mathcal{X}}_{\mathbf{a}}(t)=0$. This can be further generalized to the circumstance that $M>2$.
V-B Dynamics with drives
The drive field can be applied via the feedforward or feedback channel. Take the measurement feedback as an example, as a generalization of Section IV-A, the measurement information can be collected from an arbitrary cavity of the coupled cavity array, and the measurement information from the $m$-th cavity reads
where $x_{m}=\left(a_{m}+a^{{\dagger}}_{m}\right)/\sqrt{2}$. Then we can derive the feedback dynamics according to the feedback operator similarly as in Section IV.
For multiple coupled cavities, we take the feedback operator $G=\sum_{m}\left[\beta_{x}^{(m)}x_{m}+\beta_{p}^{(m)}p_{m}\right]$ with the feedback strength $g_{f}$, the dynamics of the operators can be generalized by Eq. (55) after averaging the homodyne detection noise as
$\left\langle\dot{\sigma}_{-}^{(m)}\right\rangle=-ig_{m}e^{i\delta_{m}t}\left\langle a_{m}\right\rangle-\mathcal{F}_{m}(t)\kappa^{(1)}\left\langle\sigma_{-}^{(m)}\right\rangle,$ (70a)
$\left\langle\dot{a}_{m}\right\rangle=-i\Delta\left\langle a_{m}\right\rangle-ig_{m}e^{-i\delta_{m}t}\left\langle\sigma_{-}^{(m)}\right\rangle-\kappa_{m}\left\langle a_{m}\right\rangle$
$-ig_{f}\left(\beta_{x}^{(m)}+i\beta_{p}^{(m)}\right)\left\langle x_{m}\right\rangle-iJ_{c}\left(\left\langle a_{m-1}\right\rangle+\left\langle a_{m+1}\right\rangle\right).$ (70b)
Obviously, the feedback drive will not directly influence the dynamics of $\left\langle\sigma_{-}^{(m)}\right\rangle$ according to Eq. (70a), but can influence the atomic dynamics via the atom-cavity couplings.